OpenStudy (anonymous):

Calculate the standard error of the mean for the sample of four: 9, 10, 9, 10. Standard deviation = .58, mean of sample. 9.5.

ganeshie8 (ganeshie8):

standard error = \(\large \frac{s}{\sqrt{n}} \)

ganeshie8 (ganeshie8):

\(\large \frac{0.58}{\sqrt{4}} \)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

i see "math processing error" on the first two that you sent

OpenStudy (anonymous):

i know its supposed to be the standard deviation divided by the square root of the sample number

OpenStudy (anonymous):

but I am not getting the right answer...it is a multiple choice question

ganeshie8 (ganeshie8):

oh ok, then latex is not working

ganeshie8 (ganeshie8):

standard error = 0.29

OpenStudy (anonymous):

i have four multiple choice options, 1.7, 1.28, .88, .62

OpenStudy (anonymous):

maybe someone made a mistake

OpenStudy (anonymous):

i will give you a medal, since I got the same answer as you anyhowwwwww

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

I emailed the professor...I guess thats all i can do!

ganeshie8 (ganeshie8):

options maybe incorrect

ganeshie8 (ganeshie8):

good luck !